OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x minus 1, divided by 2, solve for g(f(3)).

OpenStudy (math&ing001):

f(3) = 2*3 + 1 = 7 g(f(3)) = g(7) = (3*7 - 1)/2 = 10